Of the members of three athletic teams in a school 21 are in the cricket team, 26 are in the hockey team and 29 are in the football team. Among them, 14 play hockey and cricket, 15 play hockey and football, and 12 play football and cricket. Eight play all the three games. The total number of members in the three athletic teams is


1
Expert's answer
2022-01-09T15:54:55-0500

Given n(c)=21 and n(H⋂C)=14

n(H)=26 and n(H⋂F)=15

n(F)=29 and n(H⋂F⋂C)=8

∴n(C⋃H⋃F)=n(C)+n(H)+n(F)−n(H⋂C)−n(H⋂F)

−n(F⋂C)+n(H⋂F⋂C)

=21+26+29−14−15−12+8

=84−41

=43